Given U.S incomes and Expenditures as follows

wage-------- 100
Social security contributions---- 7
interest income ----- 5
personal taxes------- 20
consumption expenditure by households 60
Government purchases of goods and services 150
tranfer payments 50
coporate profit 100
invetment expenditure by business 100
corporate income tax 30
indirect business taxes 20
undistrubuted corporate profit 50
consumption of fixed capital(depreciation) 90
non income charges 40
rents 65
expenditure by foriegners 25
Net foreign factor income earned domestically 25
answer
a) using expenditure approach, what is GDP?
b)using income approach,what is GDP?
c) what is Net domestic product?
d) NI?
e) PI?
f) DI?
1
Expert's answer
2014-10-17T15:04:01-0400
A) Expenditure approach:
GDP = Y = Consumption (C) + Investment (I) + Government purchases (G) + Export (Ex) - Import (Im) = 60 + 100 + 150 + 40 = 250
B) Income approach:
GDP = Labor Income (W) + Rental Income (R) + Interest Income (i) + Profits (PR) = 100 - 20 + 65 + 5 + 100 = 250
C) NDP = GDP - Capital Consumption (CC) = 250 - 90 = 160
D) NI = GDP + NFI = 250 + 25 = 275
E) Personal income (PI) = NI - corporate taxes - retained earnings - social security + transfer payments + net interest = 275 - 30 - 50 - 7 + 50 = 238
F) DI = PI - Personal taxes = 238 - 20 = 218